OpenStudy (anonymous):

Evaluate the composition f(x)=2x-3 g(x)= x^2+2 solve (f*g)(3) Can someone pls help? Not sure how the answer is 19 from this.

OpenStudy (anonymous):

Im getting a quadratic equation when i solved this. I got 6x^3-9x^2+12x-18

OpenStudy (anonymous):

well\[(f \times g)(x)=f(x) \times g(x)\] or formally\[[f(x)][g(x)]\]

OpenStudy (anonymous):

and when i solved it i got something other than 19, so im not sure if you would want my help. lols

OpenStudy (anonymous):

thats what i did and then i multiplied that by 3 but i don't see how they got 19 at all

OpenStudy (anonymous):

(f*g) = (2x-3)(x^2+2) =2x^3+4x-3x^2-6 (f*g)(3)=2(3^3)+4*(3)-3*(3^2)-6

OpenStudy (anonymous):

you dont multiply it by 3, you plug in x=3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

hmm let me try it again real quick

OpenStudy (anonymous):

hmm i got 33

OpenStudy (anonymous):

yea, that should be it, but i guess if the book says 19, then we did something wrong.

OpenStudy (anonymous):

:( dam. lol. ugh. i literally don't see anything that can be wrong.

OpenStudy (anonymous):

sorry, couldn't be a help. i hope you can find your answer. lols

OpenStudy (anonymous):

yea, no i mean you helped me in a way. Im gona try a few other problems and see if maybe they just put the wrong answer for that problem. Have a good one.

OpenStudy (anonymous):

k, good luck

OpenStudy (anonymous):

can i ask you another question though?

OpenStudy (anonymous):

how would i do f(g(-3)) ?

OpenStudy (anonymous):

oh ok, first you would plug -3 into the g(x) equation, and then you plug whatever you got for g(x), and plug it in in f(x).

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok hold on 1 sec pls :) lol

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok ok i got 19 and it says the answer for that is 19 as well so im thinking maybe they put the wrong answer for the first one, lol.

OpenStudy (anonymous):

yay. whohoo!

OpenStudy (anonymous):

will u stay for a few more i have?

OpenStudy (anonymous):

sure, im stuck in my hw problem, just staring at it like a zombie, so why not

OpenStudy (anonymous):

lol, thanks man. maybe i can help? what is it?

OpenStudy (anonymous):

lol its calculus

OpenStudy (anonymous):

oh dam man. lol. Im in pre cal in college. Slacked off a little in high school. lol. At 1 in the morning though, its a little hard to do this lol. I haven't done this since last year.

OpenStudy (anonymous):

lol. ur est, so lets try to solve your other questions too. gotta brush my teeth and go to sleep soon. im pst

OpenStudy (anonymous):

oh ok. Yea. lol. ok hopefully im doing the other ones right but for (fxf)(-2) i got 49 and it says the answer is -17

OpenStudy (anonymous):

is it still the same function for f(x) which is 2x-3?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so its very similar to f(g(x)), except you just plug in -2 in 2x-3, which u will get -7, and plug in -7 in 2x-3, wchi is -17.

OpenStudy (anonymous):

ur just using f(x) twice. no biggie

OpenStudy (anonymous):

oh my god, lol. I just thought you took the -7 and multiplied it by -7 because thats what i thought fxf was. ugh man. Need to refresh on my math skills.

OpenStudy (anonymous):

its ok. we always get stuck. thats what learning is all about. getting stuck. hehe

OpenStudy (anonymous):

yea, lol. and thanks for the medal. umm for g(g(1/2)) am i just solving for g of 1/2 and then plugging that into g?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok i got whatever 2.25 squared plus 2 equals to.

OpenStudy (anonymous):

i suggest you put it in fractions, becuase precalc + is all about real numbers, no decimals.

OpenStudy (anonymous):

i mean 4.25 omg. srry. im sick and tiered and.. literally doing math at 2 am.

OpenStudy (anonymous):

i don't have a calc on me though. except my phone and comp. But they don't have the advanced functions.

OpenStudy (anonymous):

would that be right though?

OpenStudy (anonymous):

lemme do it on the drawing pad